OpenStudy (anonymous):

wts d integration of (x^2+a^2)dy/dx=(y+b)(x+sqrt(x^2+a^2))

OpenStudy (anonymous):

you need to solve the diff eqn?

OpenStudy (anonymous):

ya by variable separable

OpenStudy (anonymous):

x wali sari expressions ek side karlo

OpenStudy (anonymous):

its nt workin :(

OpenStudy (anonymous):

you have dy / (y+b) = (x+sqrt(x^2 + a^2))dx / (x^2 + a^2) isnt it?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so now wheres the problem?

OpenStudy (anonymous):

m nt gettin hw 2 go ahead

OpenStudy (anonymous):

ok the left hand side is clear to you isnt it??

OpenStudy (anonymous):

yup d right one s mes

OpenStudy (anonymous):

so fr the right side break the fraction into 2 parts

OpenStudy (anonymous):

\[\frac{x}{x^2 + a^2} + \frac{1}{\sqrt(x^2 + a^2)}\]

OpenStudy (anonymous):

ab aaya?

OpenStudy (anonymous):

i gt it thanks:)

OpenStudy (anonymous):

u on fb??

OpenStudy (anonymous):

can i hav ur id??

hartnn (hartnn):

O.o

OpenStudy (unklerhaukus):

\[(x^2+a^2)\frac{\text dy}{\text dx}=(y+b)(x+\sqrt{x^2+a^2})\] \[\frac{\text dy}{y+b}=\frac{x+\sqrt{x^2+a^2}}{x^2+a^2}{\text dx}\] \[\int\frac{\text dy}{y+b}=\int\frac{x+\sqrt{x^2+a^2}}{x^2+a^2}{\text dx}\]

OpenStudy (unklerhaukus):

then uses these standard integrals for \(\alpha,\beta,\gamma\in\mathbb R\) \[\int\frac{\text du}{u+\alpha}=\ln|u+\alpha|\] \[\int\frac{v}{v^2+\beta^2}\text dv=\frac12\ln|v^2+\beta^2|\] \[\int\frac1{\sqrt{w^2\pm \gamma^2}}\text dw=2\sqrt{ x\pm\gamma}=\]